# LLMs and the Human Condition

### v4

@@ 6 What LLMs are actually doing
The proposal is that the things we write down are to a very large extent descriptions of practices. Sure there are references in there to things, but LLMs don’t capture that. A Generative Predictive Transformer – or any mechanism for generative prediction – is not “understanding” that a cat has four legs, sharp teeth and fur, what it is learning is that we pat cats, we attend when they catch mice, and that they jump on window sills and knock things off. The things we notice about cats are all integrated into our own practices, and what we talk about are those practices.
What an LLM is doing to generalise from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.
7 A human super power
Although we can and do think about what other people are thinking, we generally don’t bother. Much of what we do can be attributed to either reacting to stuff around us, or to the adoption of a role in a known practice. For social animals the practices are largely shared, and visible. That is, others can see what an individual is doing, and because the practice is shared, they know what the individual will do next. A sheep dog does not need to model the mind of a sheep – the sheep’s observable actions are simply a recognisable part of a practice that sheep perform (Figure 2).
For humans, and the higher animals at least, practices can be learnt. The cat learns about the getting up in the morning routine; knows where the litter tray is and so on. People - language users - use language as part of practices, but language also empowers new practices. The symbolic nature of language reifies stuff into things so we can think about all cats when our experience is of just a few. We share these symbols - the sign and the signified - to invent (the practice of) counting and holding committee meetings, and we develop theories about "things" and how they interact. We can think about fairies, God, hermeneutics, and other people’s intentions. The proposal is that it is generally only humans that do that. As language users we inherit a comprehensive set of signs and signified with which to think. It may be the case that other animals have the ability to do that as individuals, but language provides the toolbox which is handed down to the next generation along with the practices for using it.
@@ 8 Conclusion
For the engineers wanting to build conversational user interfaces, it seems there might be potential in looking at language in use as being, not about syntax and semantics, and not a reference to things, but about negotiating a common practice between the speaker and his or her CP.

### v5

@@ 5 What LLMs are actually doing
The proposal is that the things we write down are to a very large extent descriptions of practices. Sure there are references in there to things, but LLMs don’t capture that. A Generative Pre-trained Transformer – or any mechanism for generative prediction – is not “understanding” that a cat has four legs, sharp teeth and fur, what it is learning is that we pat cats, we attend when they catch mice, and that they jump on window sills and knock things off. The things we notice about cats are all integrated into our own practices, and what we talk about is those practices.
The hypothesis is that an LLM, rather than abstracting from “cereal” to “food” and then reasoning about “hunger,” is generalising from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.
6 Summary
In large part the achievements of us humans are a product, not of individuals thinking about stuff, but of societies evolving to do what they do. A society is a set of practices that are mutually dependent and supportive; humans are the mechanism by which practices are executed. Mostly. Societies compete, and those with the better combination of practices survive. Of course modern humans can think about and change practices, and our super power is to do that with symbols, but the symbols – critically the signified referenced by the sign – needs to be identified first, and that process is grounded in the practices a society has evolved. Some practices have obvious advantage when looking at the survival of a society. Farming, child rearing and building shelters are fine, but others – going to church or a football match, elections, and street parties – are less clear and indeed it is hard to think about the consequences of doing them differently.
At all levels an individual acts a role in different practices at different times, and does not need to be aware or think about why the practice is good for anyone or anything. What the individual does need to do is to learn how to perform that role and we do that by copying. Whatever that mechanism is, it enables us to predict what happens next. This is how language works. We see a child starting the practice of getting something to eat and we all expect – including the child – that the result will be some food. When that doesn’t happen, we could reason about the contents of the mind of the mother, but we might also have learnt the practice of negotiation.
@@ 7 Conclusion
We don’t need to read minds, we just need to predict what comes next and this is what a “generative pre-trained” whatever does well when it is trained on what are primarily stories of practices being played out. For us engineers wanting to build conversational user interfaces, this looks like a useful function. It does not solve the miriad challenges for conversational machines which the promoters of the new AI seem to leave for “prompt engineers” to solve.